f(x) =lin(x+linx)


1
Expert's answer
2021-05-26T03:29:34-0400

"Given,\\f(x)=ln(x+lnx)"


and we know that "\dfrac{d}{dx}lnx=\dfrac{1}{x}"


So,

"f'(x)=\dfrac{d}{dx}ln(x+lnx)\\\ \\f'(x)=\dfrac{1}{(x+lnx)}\dfrac{d}{dx}[x+lnx]=\dfrac{1}{(1+lnx)}\cdot[1+\dfrac{1}{x}]=\dfrac{x+1}{x(1+lnx)}"


Hence, "\boxed{f'(x)=\dfrac{x+1}{x\cdot(1+lnx)}}"